SOME RATIONAL VERTEX ALGEBRAS by Dražen Adamović, Zagreb, Croatia

Abstract. Let L((n−3/2)Λ_{0}), n ∈ N, be a vertex operator algebra associated to the irreducible highest weight module L((n−3/2)Λ_{0}) for a symplectic affine Lie algebra. We find a complete set of irreducible modules for L((n−3/2)Λ_{0}) and show that every module from L((n − 3/2)Λ_{0}) from the category O is completely reducible. Mathematics subject classification: 17B67 Keywords: symplectic affine Lie algebra, vertex operator algebra, admissible weights, modules for vertex operator algebra, rational operator vertex algebra.